OpenStudy (anonymous):

The function f(x) = 396(3)x represents the growth of a dragonfly population every year in a remote swamp. Pierre wants to manipulate the formula to an equivalent form that calculates every month, not every year. Which function is correct for Pierre's purposes?

OpenStudy (anonymous):

My options are: f(x) = 33(3)^x f(x) = 396(3^12)^(1/12) f(x) = 396(3^1/12)^12x f(x) = 4752(3)^x